OpenStudy (anonymous):

Lagrange Multipliers: Find the Max and Min of x^2+y^2 given the constraint: 9x^2+4y^2=25

OpenStudy (anonymous):

I found the critical values +/- 1/3 for x and +/- 3/4 for y, but I'm not sure whether these would give me the max or min.

OpenStudy (abb0t):

Well, you know that possible solutions lies on a disk of radius 5 which is a closed and bounded region, yes.

OpenStudy (anonymous):

Not sure if my critical values are correct. Even if they are, should I just plug them back into the original function to find max and min?

OpenStudy (abb0t):

Plug it into your original fnction now that you have them.

OpenStudy (anonymous):

I tried doing that, but both coordinates gave me the same output. How would I use the constraint to find the min?

OpenStudy (abb0t):

did you take the partial?

OpenStudy (anonymous):

I took the partial of the constraint and solved for lambda: d/dx: 1=(lambda)(18x) d/dy: 1=(lambda)(8y)

OpenStudy (anonymous):

lambda=1/18x and 1/8y Not sure if that's correct.